2(3x+1)(x-7)=0
can someone please show me step-by-step how to solve this problem?
i am not understanding how to solve using the quadratic formula...

Answer by mananth(16946)

2(3x+1)(x-7)=0
(3x+1)(x-7)=0
3x(x-7)+1(x-7)=0
3x^2-21x+x-7 =0
3x^2-20x-7=0
3x^2-21x+x-7=0
3x(x-7)+1(x-7)=0
(x-7)(3x+1)=0
x=7 Or x=-1/3